# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

### 06 Feb 2024 v2

@@ 4 Related Work
We will note that while the LLM-Modulo framework is being proposed in general form here for the first time, there are certainly works in leveraging LLMs in planning and reasoning tasks that are in line with the spirit of the LLM-Modulo framework. For exmaple, both Valmeekam et al. (2023c) and Stechly et al. (2023) describe and evaluate a backprompting interaction between an LLM and an external verifier. Work on FunSearch Romera-Paredes et al. (2023) depends on a generate-test loop between a specially fine-tuned LLM that guesses solutions, and an external symbolic evaluator that critiques them. The authors note how the external verifier is critical for avoiding falling prey to hallucinations (i.e., approximate solution candidates that have flaws). AlphaGeometry Trinh et al. (2024) too depends on the Generate-Test-Critique interaction between a fine-tuned LLM and a symbolic evaluator. Both these systems fine-tune pre-trained LLMs with task specific synthetic data–the correctness of which is vetted with external simulators (as we discuss in Section 3.3).
@@
While we focused on PDDL planning tasks for the sake of concreteness, we believe that the essence of LLM-Modulo framework is equally applicable to other scenarios involving planning and reasoning–such as Reinforcement Learning with Simulators. Such RL systems rely on rewards as feedback to train a policy. Simulators takes on the roles of plan evaluation and critiques performed by the respective critics in the LLM-Modulo framework (e.g. Rajvanshi et al. (2023)). The fact that simulators play the role of verifiers is often not explicitly recognized in cases where LLMs are used as an actor to generate an admissible plan by interacting with a simulator, for example in the case of AlfWorld Yao et al. (2023b); Shinn et al. (2023) and Minecraft Wang et al. (2023). As mentioned in Section 3, similar to extracting a domain model such as in the case of PDDL planning Guan et al. (2023), designing a reward model for the plan generation - feedback cycle is yet another potential use case that has been recently looked at for text-based Kwon et al. (2022); Hao et al. (2023) and robot manipulation Ma et al. (2023) domains.
@@
Interestingly, the fact that LLM’s can help come up with approximate quasi-symbolic transition models, reward models and models of high level actions has made a bigger splash in RL. This is because for far too long, researchers there have tried to spurn any high level models (lest that would involve depending on humans; Kambhampati (2021)) and focused on learning to act from sensory information, under the name of “deep reinforcement learning.” Given the horrendous sample complexity of the DRL methods even in reaching a single subgoal, and the well known fact that even approximate symbolic models can help drastically improve the performance (c.f. Guan et al. (2022)), coupled with the fact that LLM’s are only too glad to dream up approximate models and goal recipes, there has been a performance revolution of sorts there Yao et al. (2023b); Liang et al. (2023); Wang et al. (2023). If we look beyond the improvements in these lower level goal seeking behaviors–especially in the presence of ergodic simulators, the RL approaches dependent on LLMs will encounter the same issues regarding subgoal interactions that our discussion of PDDL planning problems brought into focus. The LLM-Modulo inspired frameworks will thus, we believe, be equally relevant there. Indeed, SayCan Ahn et al. (2022) the earliest use of LLMs in generating policies in an RL-with-Simulator scenario, explicitly filters the action choices suggested by the LLM with the help of simulator.
@@
While we focused on text based LLMs (such as GPT4), recently there have also been impressive development in multi-modal LLMs (e.g. GPT4V). While multi-modality is a great addition that increases the coverage of their System 1 imagination (Figure 1), it is not clear that this gives them System 2 competence. [^12] If you know how to complete sentences, and now learned to complete dance moves, does your ability to reason/plan magically improve?
@@ 5 Conclusion
This position paper is a modest attempt to combat both over-optimism and over-pessimism about the role of LLMs in planning and reasoning tasks. Our position is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers. In support of this position, we summarized the literature questioning the claims about the planning and self-verification capabilities of LLMs by themselves. We also discussed how conflating approximate knowledge acquisition and generating executable plans of action is behind many of the claims about planning and verification abilities of LLMs. We then shared LLM-Modulo framework, our vision for a productive way to integrate the impressive idea generation/approximate knowledge provision capabilities of LLMs with external verifiers with correctness guarantees for robust and expressive planning. We discussed how planning in LLM-Modulo framework avoids inheriting the expressiveness and search-complexity limitations of traditional symbolic planners, while retaining their soundness guarantees. As we discussed, LLM-Modulo frameworks are consistent with some of the most high-profile success stories of “neuro-symbolic” architectures, including AlphaGeometry and FunSearch.
@@ Acknowledgments
The ideas discussed in this paper have evolved over a series of talks, tutorials and twitter threads. The discussions, feedback and encouragement from colleagues, including Daniel Borrajo, Tom Dietterich, Yann LeCun, Sarath Sreedharan, and Dan Weld is gratefully acknowledged.

### 12 Jun 2024 v3

@@ 5 Related Work
While the LLM-Modulo framework is being proposed in general form here for the first time, there are certainly works in leveraging LLMs in planning and reasoning tasks that are in line with the spirit of the LLM-Modulo framework. Work on FunSearch (Romera-Paredes et al., 2023) depends on a generate-test loop between a specially fine-tuned LLM that guesses solutions, and an external symbolic evaluator that critiques them. The authors note how the external verifier is critical for avoiding falling prey to hallucinations (i.e., approximate solution candidates that have flaws). AlphaGeometry (Trinh et al., 2024) too depends on the Generate-Test-Critique interaction between a fine-tuned LLM and a symbolic evaluator. Both these systems fine-tune pre-trained LLMs with task specific synthetic data–the correctness of which is vetted with external simulators.
@@
While we focused on PDDL planning tasks for the sake of concreteness, we believe that the essence of LLM-Modulo framework is equally applicable to other scenarios involving planning and reasoning–such as Reinforcement Learning with Simulators. Such RL systems rely on rewards as feedback to train a policy. Simulators takes on the roles of plan evaluation and critiques performed by the respective critics in the LLM-Modulo framework (e.g. (Rajvanshi et al., 2023)). The fact that simulators play the role of verifiers is often not explicitly recognized in cases where LLMs are used as an actor to generate an admissible plan by interacting with a simulator, for example in the case of AlfWorld (Yao et al., 2023b; Shinn et al., 2023) and Minecraft (Wang et al., 2023b). Similar to extracting a domain model such as in the case of (Guan et al., 2023), LLMs can also be used for designing a reward model or shaping the reward (Bhambri et al., 2024; Kwon et al., 2022; Hao et al., 2023; Ma et al., 2023).
@@
Interestingly, the fact that LLM’s can help come up with approximate quasi-symbolic transition models, reward models and models of high level actions has made a bigger splash in RL. This is because for far too long, researchers there have tried to spurn any high level models (lest that would involve depending on humans; (Kambhampati, 2021)) and focused on learning to act from sensory information, under the name of “deep reinforcement learning.” Given the horrendous sample complexity of the DRL methods even in reaching a single subgoal, and the well known fact that even approximate symbolic models can help drastically improve the performance (c.f. (Guan et al., 2022)), coupled with the fact that LLM’s are only too glad to dream up approximate models and goal recipes, there has been a performance revolution of sorts there (Yao et al., 2023b; Liang et al., 2023; Wang et al., 2023b). If we look beyond the improvements in these lower level goal seeking behaviors–especially in the presence of ergodic simulators, the RL approaches dependent on LLMs will encounter the same issues regarding subgoal interactions that our discussion of PDDL planning problems brought into focus. The LLM-Modulo inspired frameworks will thus, we believe, be equally relevant there. Indeed, SayCan (Ahn et al., 2022) the earliest use of LLMs in generating policies in an RL-with-Simulator scenario, explicitly filters the action choices suggested by the LLM with the help of simulator.
@@
Although we focused on text based LLMs (such as GPT4), recently there have also been impressive development in multi-modal LLMs (e.g. GPT4V). While multi-modality is a great addition that increases the coverage of their System 1 imagination (Figure 1), it is not clear that this gives them System 2 competence. [^11] If you know how to complete sentences, and now learned to complete dance moves, does your ability to reason/plan magically improve? As we discussed earlier, we can leverage VLMs for style criticism of the robot behavior (Guan et al., 2024).
Finally, our position (with published supporting evidence) that LLMs are incapable of supporting planning in autonomous modes must seem quite at odds with the current head-long rush into agentic LLMs. We believe that the latter is largely a result of confusing “acting” with “planning.” Given their ability to translate across formalisms, it is of course possible for LLMs to invoke external services–something frameworks like AutoGPT and LangChain support. But the mere ability to invoke an action doesn’t, in any way, guarantee that the course of actions thus invoked will achieve a desired state of affairs. The only way to guarantee the latter is to to support robust planning capabilities–something our LLM-Modulo frameworks strive to do.
@@ 6 Conclusion
This position paper is a modest attempt to combat both over-optimism and over-pessimism about the role of LLMs in planning and reasoning tasks. Our position is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers. In support of this position, we summarized the literature questioning the claims about the planning and self-verification capabilities of LLMs by themselves. We also discussed how conflating approximate knowledge acquisition and generating executable plans of action is behind many of the claims about planning and verification abilities of LLMs. We then shared LLM-Modulo framework, our vision for a productive way to integrate the impressive idea generation/approximate knowledge provision capabilities of LLMs with external verifiers with correctness guarantees, for robust and expressive planning. We discussed how planning in LLM-Modulo framework avoids inheriting the expressiveness and search-complexity limitations of traditional symbolic planners, while retaining their soundness guarantees. We illustrated and discussed the many roles LLMs can play in the LLM-Modulo framework. Finally, we also discussed two case studies of adapting the LLM-Modulo frameworks.
Impact Statement
This position paper takes a stance on a robust and well-founded way of leveraging Large Language Models in planning and reasoning tasks. It (i) points out the inabilities of current pre-trained LLMs to tackle planning problems, (ii) suggests some reasons as to why there are wide-spread misunderstandings about LLM planning abilities and (iii) proposes LLM-Modulo frameworks as a way to leverage LLMs for robust planning. The main consequences of realizing this position/vision is expected to be (i) sounding caution about misapplication of LLMs in autonomous modes for planning (ii) providing a way to leverage LLMs to do robust planning. Given the current interest in agentic LLMs, these insights can have significant positive impact in mission critical situations. We do not see any obvious negative societal consequences of leveraging LLMs this way (unless of course the plans are aimed at achieving malicious goals).
@@ Acknowledgments
The ideas discussed in this paper have evolved over a series of talks, tutorials and twitter threads. The discussions, feedback and encouragement from colleagues, including Sarath Sreedharan, Tom Dietterich, Yann LeCun, Daniel Borrajo, and Dan Weld is gratefully acknowledged. The adaptation of LLM-Modulo Framework for Travel Planning, discussed in Section 4 was lead by Atharva Gundawar. Kambhampati acknowledges generous support from ONR via grants N00014-18-1-2442, N14-18-1-2840 and N00014-23-1-2409, as well as gifts from J.P. Morgan, Qualcomm and Amazon.